2 - 3x < or equal to 8-x

Answer by checkley71(8403)

2-3X<=8-X
-3X+X<=8-2
-2X<=6 DIVIDING BY A NEGATIVE INTEGER CHANGES THE < TO A > SIGN.
X>=6/-2
X>=-3 ANSWER.PROOF
LET X=-2.9
2-3*-2.9<=8-(-2.9)
2+8.7<=8+2.9
10.7<=10.9